Common fixed points of Greguš type mappings

P. P. Murthy, Bhilai, India, Y. J. Cho, Jinju, Korea and Brian Fisher, Leicester, England

Keywords: Compatible mappings of type (A), weakly commuting mappings, compatible mappings, nonexpansive mappings, common fixed points AMS Subject classification: 54H25

Abstract: This paper deals with a common fixed point theorem of Greguš type under a ϕ-type contraction. Our theorem also improves results of Fisher and Sessa [2] and Sessa and Fisher [10].